# A reduced-order framework for temperature estimation in food freezing from optimally located sensors, including turbulent conjugate flow scenarios

### v1

@@ 5.3 Temperature reconstruction with an optimal sensor placement algorithm
In this section, we test the optimal sensor placement algorithm proposed in section 4.3 using the same set of parameters for the boundary conditions used in Section 5.2. We compare the reconstruction output of this algorithm with a regularly chosen observation region, as depicted in Figure 16.a. Nested squared observation regions are taken, excluding the food region, leading to 4 sensors set-ups, namely with 56, 162, 266 and 352 (full domain excluding food) measurements. In contrast, a sparser sensor selection is made when using the 3-step algorithms, as shown in Figure 16. Interestingly, the method tends to privilege both the food surroundings and the upper part of the computational domain, corresponding to the region where the turbulent airflow interacts with the food and with the stratified air in the upper part of the cabinet.
(a) Regular placement
(b) Greedy placement (picked sensors in gray)
Figure 16: Optimal sensor placement comparison with regular squared measured regions. The Greedy algorithm prioritizes the ROM observability, thus enabling a better stability of the inverse problem.
@@
To evaluate and compare the performance of the proposed Greedy algorithm against regular sensor placement, we compute the relative l2-norm error of temperature defined in Eq. (15). The time-dependent and time-averaged values of the relative l2-norm error between regular and optimal sensor placements with different amount of measurements are compared in Figure 17 a), and 17 b), respectively.
@@
The results demonstrates the robustness of the Greedy algorithms also when decreasing the number of measurements. We observe how properly placing the sensors improves the quality of the reconstructions during the entire freezing process, achieving errors close to 1% over the whole time-dependent dynamics. Figure 17.a shows that the optimal placement of 162 sensors achieves a similar error (in $\ell^{2}$-norm) over time as the regular placement with almost twice as many sensors ($m$ = 352). Moreover, when comparing the time-averaged errors for a given number of sensors (Figure 17.b), the Greedy algorithm outperforms the regular placement, especially for a small number of measurements (such as for $m=52$), reaching similar time-average errors that those obtained with full-domain measurements (thermocamera).
(a) Time-dependent $\ell^{2}$-norm error.
(b) Time-averaged $\ell^{2}$-norm error.
Figure 17: Comparison of the reconstruction quality of the regular sensor placement and Greedy algorithm for optimal sensor placement. The Greedy algorithm allows for considerable reduction of the number of sensors without any significant error increase.
@@
Similar to the previous sections, we evaluate the reconstruction quality for two different local quantities of interest. This time, we compare the ground-truth with the Greedy algorithm for the three set of sensors, i.e. $GA_{1}$ for $m=52$, $GA_{2}$ for $m=162$, and $GA_{3}$ for $m=266$, for the local evolution of temperature in Figure 18 a), and for the local freezing rate in Figure 18 b). Additionally, we compute the accumulative relative error over time by integrating in time the Equation 16.
@@
From Figure 18 a) and b), we can highlight that the local temperature evolution and the freezing rate are well reproduced with the Greedy strategy using $GA_{1}$ and $GA_{3}$ in points P1 and P3. For point P2, $GA_{3}$ depicts a better reconstruction quality, particularly in the freezing rate during the phase-changing stage. Regarding the accumulative relative error over time, we can highlight that half of the total temperature error and two-thirds of the total freezing rate error accumulated at the end of the process comes from the initial chaotic stage, which is associated with the unsteady turbulent fluid dynamics and the rapid cooling of the food boundaries. At this stage, the temporal derivative of the reconstructed field is of relevance. Despite this, the method reconstructs the temperature field very well and even achieves a good definition of its temporal evolution, which is not trivial since the proposed algorithm does not incorporate any condition regarding the temporal derivative of the reconstructed field.
(a) Local evolution of temperatures and their accumulative relative errors.
(b) Local evolution of freezing rates and their accumulative relative errors.
Figure 18: Comparison of local quantities for three fixed points inside the food between the ground-truth and the reconstruction with the Greedy algorithm using three sets of measurements.
@@ 6 Conclusions
This study addresses the inverse estimation of temperature fields in food freezing processes, a critical factor in ensuring product quality in the food industry. Numerical simulations of thermally-coupled flow systems, including liquid-to-solid phase change and turbulence modeling, have been used to generate a synthetic dataset and to create a physics-informed reduced-dimensional space to regularize the optimization problem for the state estimation . The approach allowed to effectively reconstruct the time-dependent temperature distribution within the food matrix from sparse sensor data, even when observations inside the food were completely excluded from the available dataset. The optimal placement of temperature sensors was determined using a new greedy algorithm based on available a-priori estimates on the reconstruction, enhancing estimation accuracy and minimizing the number of sensors required. The novelty of using this approach is two-fold. Firstly, this strategy has not yet been employed in the context of food freezing process. Secondly, the underlying mechanism of the Greedy approach is based on a theoretical work on a-priori bounds (14), and, unlike more traditional sensor placement approaches, it does not directly relies on online phase information. In other words, the strategy only requires the ROM and the position of the sensors, but not the values of the measurements themselves, allowing for general conclusions independent of the data being assimilated at a given iteration. The Greedy approach has been successfully validated for defining sensor locations that maximized data representativeness, offering a cost-effective solution for complex temperature monitoring applications in food preservation.
Future work will explore further refinements in the reconstruction algorithm, including model-bias [38], and/or measurement-bias [43], control over the time derivative of the reconstructed field, as well as real-time estimation techniques to advance food safety protocols.
@@ Acknowledgments
E. Castillo acknowledges the support provided by ANID-Chile throughout the project FONDECYT 1210156. D.R.Q. Pacheco acknowledges funding by the Federal Ministry of Education and Research (BMBF) and the Ministry of Culture and Science of the German State of North Rhine-Westphalia (MKW) under the Excellence Strategy of the Federal Government and the Länder. F. Galarce acknowledges the support provided by DI-VINCI PUCV funding 039.482/2024. Diego. Rivera acknowledges the support granted by the DICYT-POSTDOC 052416CD funded by VRIDEI-USACH.

### v2

@@ 5.4 Temperature reconstruction with an optimal sensor placement algorithm
In this section, we test the optimal sensor placement algorithm proposed in Section 4.3 using the same set of parameters for the boundary conditions used in Section 5.3. We compare the reconstruction output of this algorithm with a regularly chosen observation region, as depicted in Fig. 15. Nested squared observation regions are taken, excluding the food region, leading to 4 sensor set-ups, namely with 56, 162, 266, and 352 (full domain excluding food) measurements. In contrast, a sparser sensor selection is made when using the 3-step algorithms, as shown in Fig. 14. Interestingly, the method tends to privilege both the food surroundings and the upper part of the computational domain, corresponding to the region where the turbulent airflow interacts with the food and with the stratified air in the upper part of the cabinet.
Figure 14: Greedy sensor placement. We compare the temperature reconstruction using a fixed number of measurements ($m_{1}=56$, $m_{2}=162$, $m_{3}=266$ and $m_{4}=352$), always discarding the food chunk. We observe how the methodology provides a sparse spatial sampling of the domain based on the pre-built reduced order model.
@@
To evaluate and compare the performance of the proposed Greedy algorithm against regular sensor placement, we compute the relative l2-norm error of temperature defined in Eq. (15). The time-dependent and time-averaged values of the relative l2-norm error between regular and optimal sensor placements with different amount of measurements are compared in Fig. 15 $(b)$, and 15$(c)$, respectively.
@@
The results demonstrates the robustness of the Greedy algorithms also when decreasing the number of measurements. We observe how properly placing the sensors improves the quality of the reconstructions during the entire freezing process, achieving errors close to 1% over the whole time-dependent dynamics. Fig. 15$(a)$ shows that the optimal placement of 162 sensors achieves a similar error (in $\ell^{2}$-norm) over time as the regular placement with almost twice as many sensors ($m$ = 352). Moreover, when comparing the time-averaged errors for a given number of sensors (Fig. 15, the Greedy algorithm outperforms the regular placement, especially for a small number of measurements (such as for $m=52$), reaching similar time-average errors that those obtained with full-domain measurements (thermo-camera). We observe a good behavior against the number of sensors in Fig. 15$(b)$, where the method manages to keep the error low despite the diminished amount of measurements and we also observer a good sensitivity for different freezing conditions in Fig. 15$(c)$, where several boundary conditions and physical parameters are tested out, delivering an average error in time no large than 2 %.
Figure 15: Numerical study with greedy algorithm for sensor placement. We observe how the method selects sensors to ensure better accuracy in (a). The method delivers a good behavior against the number of measurements, showing much better convergence rate in (b) with respect to the regular sensor grid selection. In (c), we see this result systematically tested against 16 different simulations, showing good sensitivity against a number of freezing configurations including different boundary conditions.
@@
Similar to the previous sections, we evaluate the reconstruction quality for two different local quantities of interest. This time, we compare the ground-truth with the Greedy algorithm for the three set of sensors, i.e. $GA_{1}$ for $m=52$, $GA_{2}$ for $m=162$, and $GA_{3}$ for $m=266$, for the local evolution of temperature in Fig. 16$(a)$, and for the local freezing rate in Fig. 16$(b)$. Additionally, we compute the accumulative relative error over time by integrating in time the Equation 16. Additionally, we should note that the sensor location is magnitude-independent, meaning that we only need as an input a set of location candidates, making the greedy search only needed once for all the test cases.
@@
From Fig. 16$(a)$ and $(b)$, we can highlight that the local temperature evolution and the freezing rate are well reproduced with the Greedy strategy using $GA_{1}$ and $GA_{3}$ in points P1 and P3. For point P2, $GA_{3}$ depicts a better reconstruction quality, particularly in the freezing rate during the phase-changing stage. Regarding the accumulative relative error over time, we can highlight that half of the total temperature error and two-thirds of the total freezing rate error accumulated at the end of the process comes from the initial chaotic stage, which is associated with the unsteady turbulent fluid dynamics and the rapid cooling of the food boundaries. At this stage, the time derivative of the reconstructed field is of relevance. Despite this, the method reconstructs the temperature field very well and even achieves a good definition of its temporal evolution, which is not trivial since the proposed algorithm does not incorporate any condition regarding the time derivative of the reconstructed field.
Figure 16: Temperature evolution (a), freezing rates (b), and cumulative errors (c,d), for the food chunk in 3 control points using the greedy reconstruction algorithm. We observe, as expected, an excellent behavior for the temperature reconstruction within the food piece, whereas the freezing rate exhibits an admittedly larger error, as no control is made over the field derivatives during the reconstruction process.
@@ 6 Conclusions
This study has addressed the inverse estimation of temperature fields in food-freezing processes, a critical factor in ensuring product quality in the food industry. Numerical simulations of thermally coupled flow systems, including liquid-to-solid phase change and turbulence modeling, have been used to generate a synthetic dataset and create a physics-informed reduced-dimension space to regularize the optimization problem for the state estimation. The approach effectively estimates the time-dependent temperature distribution inside the food from sparse sensor data, even when observations exclude the food domain.
The optimal placement of temperature sensors used for reconstruction was determined using a new greedy algorithm based on available a priori estimates, which enhances accuracy and minimizes the number of sensors required. Furthermore, we designed a real-time reconstruction approach, which involves solving a linear system of order $\sim 10^{2}$, applicable to any new dataset. The novelty of using this approach goes beyond the fact that this strategy had not yet been employed in food freezing applications; most importantly, our approach only requires the ROM and the sensor positions, but not the values of the measurements themselves, enabling the sensor location computation a-priori and independent of the data being assimilated at a given iteration. The Greedy approach has been successfully validated for defining sensor locations that maximized data representativeness, offering a cost-effective solution for complex temperature monitoring applications in food preservation.
Future work will explore further refinements in the reconstruction algorithm, including model bias [35], and/or measurement bias [57], control over the time derivative of the reconstructed field, as well as real-time estimation techniques to advance food safety protocols. From the physical viewpoint, future extensions will include Large-Eddy Simulation turbulent models instead of URANS strategies and the temperature estimation of liquid-to-solid scenarios where the natural convection inside the freezing food cannot be despised. The application of the proposed framework for multiphase food freezing or 3D simulations is also a path to strengthen the credibility of the model in the food industry.
@@ Acknowledgments
E. Castillo acknowledges the support provided by ANID-Chile throughout the project FONDECYT 1210156. D.R.Q. Pacheco acknowledges funding by the Federal Ministry of Education and Research (BMBF) and the Ministry of Culture and Science of the German State of North Rhine-Westphalia (MKW) under the Excellence Strategy of the Federal Government and the Länder. F. Galarce acknowledges the support provided by DI-VINCI PUCV funding 039.731/2025. Diego. Rivera acknowledges the support granted by the DICYT-POSTDOC 052416CD funded by VRIDEI-USACH.